OpenStudy (anonymous):

Find the polynomial f(x) that has the roots of -3, 5 of multiplicity 2. Explain how you would verify the zeros of f(x)

OpenStudy (anonymous):

It is easy! ¿What are the roots? They are the values of X that makes f ( x ) = 0. Then, you will take advantage of the fact that if one factor of a product becomes cero, then all the products become cero. For example (a - x)(b + x) If the first (a - x) or the second (b + x) factor becomes cero, all the product is cero, is it ok?

OpenStudy (anonymous):

So you wouldhave (x+3) (x-5)

OpenStudy (anonymous):

Then...you could use that knowledge to build your function. You want a root of x = -3 ? Ok, then we need a factor that turns into cero when x = -3. It is (x + 3) , so x = -3 , then (-3+3) = 0. You do the same with the other root x = 5, so...(x+3)(x-5)

OpenStudy (anonymous):

Yes you do

OpenStudy (anonymous):

Ok what about the multiplicty partof it

OpenStudy (anonymous):

Is redundant...multiplicity 2 is to have a x^2 as the "max" exponent. If you distribute that you will arrive at some expression of the form ax^2 + bx + c and that is the form of all the polinomics of multiplicity two, is just a quadratic.

OpenStudy (anonymous):

So what would be the polynomial im so confused lol

OpenStudy (anonymous):

The polinomial is what you have, (x+3)(x-5) but you have to do the distribute. x^2 -5x + 3x -15 = x^2 -2x - 15 Then you have a quadratic equatin. And a quadratic equation is a polinomial of multiplicity 2.

OpenStudy (anonymous):

A quadratic equation is a KIND of polinomial of multiplicity 2.

OpenStudy (anonymous):

x^3 + x +1 And this is a polinomial of multiplicity 3. (The higher exponent is what matters)

OpenStudy (anonymous):

Ok thanks and its says how would you verify the zeros is that through synthetic division?

OpenStudy (anonymous):

Read the first three paragraphs; http://www.purplemath.com/modules/synthdiv.htm

OpenStudy (anonymous):

I don´t know exactly what are you meaning with "synthetic division" because I learned maths in spanish and some words or "titles of the concepts" escape me :P. I guess the first three paragraphs will help us.

OpenStudy (anonymous):

how would you check if the zeros are right

OpenStudy (anonymous):

Give X the values x = 5 and x = -3

OpenStudy (anonymous):

would you subsitute it into what the equation was that you got

OpenStudy (anonymous):

But is has no sense...because you have just build the expression so that 5 and -3 are roots!

OpenStudy (anonymous):

f(x) = (x-5)(x+3) or which is the same f(x) = x^2 -2x - 15 Let x be -3 and compute, then do the same with 5.

OpenStudy (anonymous):

Did I convince you? Is it ok?

OpenStudy (precal):